OpenStudy (anonymous):

Sally initially has 4 hours of pop music and 3 hours of classical music in her collection. Every month onwards, the hours of pop music in her collection is 8% more than what she had the previous month. Her classical music does not change. Which function shows the total hours of music she will have in her collection after x months?

OpenStudy (anonymous):

f(x) = 4(1.08)x + 3 f(x) = 3(1.08)x + 4 f(x) = 4(0.08)x + 3 f(x) = 3(0.08)x + 4

OpenStudy (anonymous):

GET A LIFE

OpenStudy (anonymous):

Help

OpenStudy (anonymous):

Btw I have more of a life than you

OpenStudy (anonymous):

please help @poopsiedoodle

OpenStudy (anonymous):

Do you understand this?

OpenStudy (anonymous):

Would it be C? @poopsiedoodle

OpenStudy (poopsiedoodle):

I haven't the slightest idea. I'm looking it up on teh google.

OpenStudy (poopsiedoodle):

holy crap google isn't loading what even

OpenStudy (anonymous):

I say C or D

OpenStudy (anonymous):

From what the gave in the problem

OpenStudy (poopsiedoodle):

sorry, my net went out. It's 4, since pop is the one that's changing.

OpenStudy (poopsiedoodle):

oops. I meant C.